Calculate the ΔH for the reaction: CS2 (l) + 2 O2 (g) 🡺 CO2 (g) + 2 SO2 (g)


1
Expert's answer
2021-06-18T07:32:33-0400

"\Delta{H}_{reaction}=\sum{\Delta{H_{products}-\sum\Delta{H}_{reactants}}}=-393.5+2(-296.8)-87.9=-1075kJ/mol"


Answer: -1075 kJ/mol